Saturday 02/11/12

11:00A — 11:45 P

  • Collect hard disk and slide from O/N STORM run
  • analyzing confocal data from S44 (antp/ubx/H3K4).
  • new confocal images: Pc, Su(Z)12, S44
  • making new antibodies (with KE) see records:
  • 12P remove samples from oven (13 hours) let cool.
  • 3:45P test samples — still soft.  sample cracks when squeezed like very thin ice over a gel.   Previously observed resin component B much fainter color than previous batches.
  • returned sample to oven at 70C (4P).
  • New confocal of S44 (antp/ubx/H3K4) and S45 (13 intron), previous data not saved well.
  • STORM Psc (2 color).  O/N STORM of S45 13 intron + H3K4 (4 color)
This entry was posted in Summaries. Bookmark the permalink.